If the direct quote for the rand-dollar exchange rate is R19,07 = $1,00, what will the
indirect quote be?
1
Expert's answer
2021-07-16T09:36:33-0400

"Indirect \space quote=\frac{1}{direct\space quote}"

direct quote"=\frac{19.07}{1.00}=19.07"

Indirect quote"=\frac{1}{19.07}=0.052438384"


$1.00=R0.052438384